Let A, B, C be three sets then A U B = C, A revU B = phi prove a=c-B 

Grade:11

1 Answers

Arun
6 years ago
As we know n(A)+n(B) = n(A U B)+ n(A revU B)
hence put the values which are given
A+B= C+ 0
then A=C –B
hence proved